OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL!! What is the slope of the line that passes through the points (1, 3) and (5, –2)? 5/4 -5/4 4/5 -4/5

jigglypuff314 (jigglypuff314):

\[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given the points (x1, y1) and (x2, y2) so to find the slope between (1, 3) and (5, –2) plug in x1=1 , y1=3 , x2=5 , y2=-2

OpenStudy (anonymous):

tbh i dont understand ??

jigglypuff314 (jigglypuff314):

when you plug it all in you would get \[\frac{ (-2)-(3) }{ (5)-(1) }\]

OpenStudy (anonymous):

ok